# SLGaussian: Fast Language Gaussian Splatting in Sparse Views

### v1

@@ Abstract
1 Introduction
@@
While NeRF-based methods [12, 21, 47] and 3DGS-baesd mothods [33, 35, 40, 19] have made significant advances in 3D language scene learning, they rely heavily on dense multi-view information, which makes them unsuitable for scenarios with limited viewpoints. When only a few views are available, it becomes challenging to reconstruct detailed 3D geometry and maintain consistent language information across those views. This hinders the system’s ability to accurately perform 3D language field learning, as the semantic information cannot be reliably integrated across different perspectives. Additionally, existing methods often require per-scene optimization, which is not only time-consuming but impractical for real-world applications, such as robotics, autonomous navigation, or industrial monitoring, where rapid inference is necessary. In such cases, the inability to perform direct inference on 3D scenes limits their utility in providing efficient semantic understanding of the environment.
Figure 1: With just two RGB views, our method infers 3D semantic fields in under 30 seconds without per-scene optimization. On LERF and 3D-OVS datasets (image resolution 416 × 576), query time is 0.011 seconds/query, outperforming existing methods in both speed and IoU metrics.
@@
To overcome these limitations and enable the broader application of semantic fields in real-world scenarios, we focus on improving inference speed and open-vocabulary query quality through three key innovations. First, inspired by recent feed-forward approaches, we design a feed forward method that allows for direct inference without the need for per-scene optimization after model convergence, greatly reducing time costs, which is based on 3DGS [20] for its fast and photorealistic rendering to accelerate the rendering process further. Additionally, to enable open-vocabulary query, related works learn high-dimensional CLIP [34] features or use additional networks to reduce the dimensionality of these features, which often hampers inference speed. We will seek a more efficient feature embedding method to overcome this limitation and improve query performance. Moreover, the existing methods typically do not address the issue of semantic feature inconsistency caused by differing viewpoints, which can lead to a degradation in the quality of reconstruction. We will explore a method to resolve the problem of semantic feature inconsistency.
@@
Based on these insights, we present the first feed-forward method for constructing semantic fields from sparse viewpoints. Our model takes sparse-view RGB images as input and utilizes two branches: one for predicting base Gaussian parameters and another for semantic parameters, enabling fast construction of 3DGS-based semantic fields. For base Gaussian parameters, we build on existing 3DGS feed-forward methods like [3, 6], while a dedicated module handles feature extraction and mapping for semantic predictions. After predicting both sets of parameters, we render 3D semantic features using the splatting technique. Then we introduce a multi-view language memory bank with low-dimensional semantic label IDs, which also facilitates efficient matching between rendered semantic features and query prompts. Additionally, we address inconsistent multi-view segmentation from the SAM model through our mask association method. In sparse-view experiments on the LERF and 3D-OVS datasets, our method achieves sharp 3D object segmentation boundaries and outperforms existing approaches in chosen IoU, Localization Accuracy, and mIoU. At a resolution of 416 × 576, our method delivers scene inference in around 25 seconds and query response times of 0.011 seconds, significantly faster than current methods.
@@ 2 Related Work
NeRF and 3D Gaussian Splatting (3DGS). In recent years, Neural Radiance Fields (NeRF) [31] and 3D Gaussian Splatting (3DGS) [20] have made significant advancements in 3D scene modeling and rendering. NeRF [31] synthesizes photorealistic scenes through volumetric rendering and multi-view consistency backpropagation, but its reliance on ray marching results in slow rendering speeds. Subsequent research, such as mip-NeRF [1] and TensoRF [4], improved its rendering efficiency and multi-resolution representation. Unlike NeRF, 3DGS employs explicit 3D Gaussian distributions instead of MLPs, achieving real-time and efficient rendering through adaptive density control and tile-based rasterization. Tang et al. [37], Liu et al. [27] demonstrated the potential of 3DGS in generative tasks. Methods like [44, 39, 25, 43, 15] enhanced Gaussians’ flexibility in handling dynamic and complex scenes by learning 3D Gaussians in canonical space combined with deformation fields. Zhang et al. [50] introduces separated intrinsic and dynamic appearance feature to capture the unchanged scene appearance along with dynamic variation like illumination and weather. Additionally, methods like [33, 40, 35] incorporate language information into 3DGS scene representation, enabling high-quality open-vocabulary 3D scene queries.
@@
Sparse View Reconstruction. Although techniques of NeRF and 3DGS perform well in dense-view scenarios, they face challenges in sparse-view settings, particularly in maintaining high-quality scene generation under sparse data conditions, necessitating further optimizations to improve generation quality and efficiency. Methods like [48, 3, 6] use a feed-forward manner to learn the 3D scene, which take images as input and predict the parameters of the 3D representation. Specifically, PixelNeRF [48] uses convolutional neural networks to extract features from input contexts, addressing the challenges of low-quality reconstructions caused by sparse views. PixelSplat leverages epipolar transformers to extract scene features, while MVSplat uses cost volumes to improve scene feature extraction. FreeNeRF further improves reconstruction by employing frequency and density regularization strategies to reduce artifacts from insufficient inputs, without additional computational costs. To mitigate overfitting in 3DGS [20] models under sparse views, approaches like FSGS [53] and SparseGS [41] introduce depth estimators to regularize the optimization process. FreeSplat [38] builds on this by incorporating not only cost volume construction but also multi-scale feature aggregation, allowing accurate 3D Gaussian distribution localization and supporting free-viewpoint synthesis. These advancements demonstrate the growing potential of sparse-view reconstruction technologies. Learning from the 3DGS-based generative frameworks [3, 6, 38, 49, 42, 7, 45], SLGaussian adopts a similar approach to infer high-quality 3DGS semantic scenes in sparse-view semantic learning tasks.
@@
3D Semantic Field. Integrating semantic information into 3D scenes is crucial for downstream fields such as robotics, AR, and VR. 2D semantic segmentation mathods like [22] and 2D open-vocabulary segmentation methods like [24, 11, 26] have inspired further exploration in understanding 3D scenes. Zhi et al. [51] first introduced semantic layers in NeRF, laying the foundation for deeper scene understanding. Subsequent studies, such as LERF [21], ISRF [12], and DFF [47], have incorporated more refined semantic segmentation techniques, significantly improving the level of detail in 3D visual representations. Furthermore, 3DGS technology has also made breakthroughs in semantic tasks. The foundational work of [20] and the subsequent developments of EgoLifter [13], SAGA [2], Gaussian Grouping [46], and Gaga [29] have greatly optimized the application of 3DGS in semantic segmentation. LangSplat[33], LEGaussians[35], OpenGaussian[40] and FastLGS [19] further explored the integration of language information into 3DGS, deepening semantic interpretation in 3D scene modeling. More previous methods like [5, 9, 14, 17, 18, 23, 30, 32, 36] have also explored embedding open-vocabulary information into 3D scenes to enable open-vocabulary interaction. However, these techniques often exhibit limitations in handling sparse view inputs, especially when viewpoints are highly limited, making it difficult to effectively manage semantic consistency between views. The lack of sufficient RGB information leads to suboptimal reconstruction quality of both 3D scenes and their semantic content, highlighting the urgent need for efficient and accurate 3D semantic learning under sparse views.
@@
Figure 2: Overview of our approach. Starting with two-view RGB images, we first apply SAM [22] and a video object tracking model to generate consistent labeled masks (Associated Masks) and construct the language memory bank to store CLIP embeddings [34] for each label ID (Section 3.2). The RGB images and associated masks are processed by the Base Gaussians Prediction and Semantic Parameters Prediction modules, which are then combined to generate semantic Gaussians. These are then used to render the 3D semantic field (Section 3.1). Finally, we perform Open-vocabulary Querying through our Multi-view Language Memory Bank (Section 3.2 and Section 3.3), enabling efficient querying of objects within the 3D scene.
3 Proposed Approach
@@
3.1 3DGS-based Semantic Parameters Prediction
@@
Base Gaussians Prediction. Inferring the semantic field directly from RGB images makes it difficult for the model to fully capture the scene’s geometric structure. A robust semantic field representation requires a deep understanding of the scene’s geometry. To address this, we first apply 3D scene reconstruction methods to provide the model with a foundational grasp of the scene, which then guides the learning of semantic features. As a result, we construct a regular color field using a feed-forward 3D Gaussian Splatting (3DGS) [20] method to derive the base Gaussians:
$\{{c}_{i},{o}_{i},{\mu}_{i},{\delta}_{i}\}=\text{GS-Pred}\left\{\{R_{j}^{3\times H\times W}\}\right\},$ (1)
@@
here, ${c}_{i}$, ${o}_{i}$, ${\mu}_{i}$, and ${\delta}_{i}$ represent the color, opacity, mean, and covariance of the $i$-th Gaussian, respectively. The term $R_{j}^{3\times H\times W}$ refers to the RGB image from the $j$-th view, where $H$ and $W$ correspond to the height and width of the image. It is important to note that in this module, we have the flexibility to use any feed-forward model. In our work, we utilize an established method [6] to predict the base Gaussians. These base Gaussian parameters are essential for accurately representing both the scene’s geometry and appearance, directly affecting the quality of semantic field reconstruction. The number of Gaussian splats in this method is determined by the input views and the image resolution, ensuring that each pixel is associated with a corresponding Gaussian point. As a result, the prediction of base Gaussians plays a crucial role in structuring the scene representation. We can either train a complete model ourselves or use pre-trained model weights from existing approaches. Once trained, the feed-forward model in this module is frozen to maintain stability for subsequent semantic field learning.
@@
Semantic Parameters Prediction. Besides predicting the base Gaussian parameters, we also predict the 3DGS semantic parameters in another branch. First, we process the input RGB images from different viewpoints to obtain consistent 2D segmentation maps across views, denoted as $\{S_{i}^{1\times H\times W}\}$, where $H$ and $W$ represent the height and width of the image, respectively. Then, during the construction of the language memory bank, which will be detailed later, we map the segmentation maps into semantic label images $\{M_{i}^{3\times H\times W}\}$. The dimensional extension to 3 is necessary because a 1D index space cannot effectively separate the semantic labels representing different objects within the scene. Next, we apply a simple CNN network to extract features from $M_{i}$. After being input into the network, $M_{i}$ first passes through a convolutional layer and then through multiple downsampling operations based on Residual Blocks [16]. Finally, it is upsampled back to the original resolution and added to the original input $M_{i}$. The resulting feature map is flattened and then passed through an MLP, which maps it to the Gaussian semantic feature parameters $f_{i}$. These parameters are 3-dimensional, corresponding directly to the 3D labels of $M_{i}$. Finally, $f_{i}$ is combined with the previously obtained base Gaussian parameters $o_{i},\mu_{i},\delta_{i}$ to generate the final 3D semantic field representation $\{o_{i},\mu_{i},\delta_{i},f_{i}\}$.
@@
$G(x)=\exp\left(-\frac{1}{2}(x-\mu)^{\top}\Sigma^{-1}(x-\mu)\right),$ (2)
@@
$F(v)=\sum_{i\in\mathcal{N}}f_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (3)
here, $C(v)$ represents the rendered color at pixel $v$ for a specific view, and $\mathcal{N}$ denotes the number of Gaussians in the tile. $\alpha_{i}=o_{i}G_{i}^{2D}(v)$, where $G_{i}^{2D}(\cdot)$ refers to the projection of the $i$-th Gaussian onto the 2D image plane.
3.2 Multi-view Language Memory Bank
Mask Association. Directly using the multi-view segmentation results inferred by the SAM model as input for our model can lead to confusion, as the inconsistent 2D segmentations across views interfere with the spatial mapping of semantic labels, as shown in our supplementary material. Existing methods mitigate this issue by using dense viewpoint information and multi-scale segmentation of SAM results. However, when reduced to the extreme sparse scenario, especially when only two views are available, this multi-scale approach to SAM segmentation becomes ineffective. Research on Gaussian scene segmentation [46, 10] treats multi-view images as consecutive frames in a video and applies a video object tracking technique [8] to unify the multi-view 2D segmentation masks inferred by SAM, achieving relatively good results. We adopted this method to enforce multi-view consistency on SAM’s inference results. However, when only two sparse-view images are directly processed by the video tracking model, the lack of sufficient reference views often result in poor segmentation consistency. To ensure more stable processing results for sparse-view image inputs, we replicated the input images five times to create a 10-frame sequence and processed it using the model, resulting in a set of 10 segmentation images with matched indices: $\{T_{11},T_{12},\dots,T_{15},T_{21},T_{22},\dots,T_{25}\}$, where $T_{ij}$ represents the $j$-th frame segmentation corresponding to the $i$-th view. Each pixel in the segmentation images is labeled with the corresponding object index. For each pixel $v$ in the $i$-th view, we choose the category through a voting algorithm:
$S_{i}(v)=\operatorname*{argmax}_{x\in\{T_{ij}(v)\}}\text{count}(x),$ (4)

### v3

@@ Abstract.
3D Gaussian Splatting, Semantic field, Open-vocabulary query, Feed-forward, Sparse-view reconstruction
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3754964isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Computer visionccs: Computing methodologies Rendering
Figure 1. With just two RGB views, our method infers 3D semantic fields in under 30 seconds without per-scene optimization. On LERF and 3D-OVS datasets (image resolution 416 × 576), query time is 0.011 seconds/query, outperforming existing methods in both speed and IoU metrics.
1. Introduction
@@
While NeRF-based methods (Goel et al., 2023; Kerr et al., 2023; Yen-Chen et al., 2022) and 3DGS-baesd mothods (Qin et al., 2024; Shi et al., 2024; Wu et al., 2024a; Ji et al., 2024; Li et al., 2025b, a, 2024b) have made significant advances in 3D language scene learning, they rely heavily on dense multi-view information, which makes them unsuitable for scenarios with limited viewpoints. When only a few views are available, it becomes challenging to reconstruct detailed 3D geometry and maintain consistent language information across those views. This hinders the system’s ability to accurately perform 3D language field learning, as the semantic information cannot be reliably integrated across different perspectives. Additionally, existing methods often require per-scene optimization, which is not only time-consuming but impractical for real-world applications, such as robotics, autonomous navigation, or industrial monitoring, where rapid inference is necessary. In such cases, the inability to perform direct inference on 3D scenes limits their utility in providing efficient semantic understanding of the environment.
@@
To overcome these limitations and enable the broader application of semantic fields in real-world scenarios, we focus on improving inference speed and open-vocabulary query quality through three key innovations. First, inspired by recent feed-forward approaches, we design a feed forward method that allows for direct inference without the need for per-scene optimization after model convergence, greatly reducing time costs, which is based on 3DGS (Kerbl et al., 2023) for its fast and photorealistic rendering to accelerate the rendering process further. Additionally, to enable open-vocabulary query, related works learn high-dimensional CLIP (Radford et al., 2021) features or use additional networks to reduce the dimensionality of these features, which often hampers inference speed. We will seek a more efficient feature embedding method to overcome this limitation and improve query performance. Moreover, the existing methods typically do not address the issue of semantic feature inconsistency caused by differing viewpoints, which can lead to a degradation in the quality of reconstruction. We will explore a method to resolve the problem of semantic feature inconsistency.
@@
Based on these insights, we present the first feed-forward method for constructing semantic fields from sparse viewpoints. Our model takes sparse-view RGB images as input and utilizes two branches: one for predicting base Gaussian parameters and another for semantic parameters, enabling fast construction of 3DGS-based semantic fields. For base Gaussian parameters, we build on existing 3DGS feed-forward methods like (Charatan et al., 2024; Chen et al., 2024a), while a dedicated module handles feature extraction and mapping for semantic predictions. After predicting both sets of parameters, we render 3D semantic features using the splatting technique. Then we introduce a multi-view language memory bank with low-dimensional semantic label IDs, which also facilitates efficient matching between rendered semantic features and query prompts. Additionally, we address inconsistent multi-view segmentation from the SAM model through our mask association method. In sparse-view experiments on the LERF and 3D-OVS datasets, our method achieves sharp 3D object segmentation boundaries and outperforms existing approaches in chosen IoU, Localization Accuracy, and mIoU. At a resolution of 416 × 576, our method delivers scene inference in around 25 seconds and query response times of 0.011 seconds, significantly faster than current methods.
@@ 2. Related Work
NeRF and 3D Gaussian Splatting (3DGS). In recent years, Neural Radiance Fields (NeRF) (Mildenhall et al., 2021) and 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) have made significant advancements in 3D scene modeling and rendering. NeRF (Mildenhall et al., 2021) synthesizes photorealistic scenes through volumetric rendering and multi-view consistency backpropagation, but its reliance on ray marching results in slow rendering speeds. Subsequent research, such as mip-NeRF (Barron et al., 2021) and TensoRF (Chen et al., 2022), improved its rendering efficiency and multi-resolution representation. Unlike NeRF, 3DGS employs explicit 3D Gaussian distributions instead of MLPs, achieving real-time and efficient rendering through adaptive density control and tile-based rasterization. Tang et al. (2023); Liu et al. (2024) demonstrated the potential of 3DGS in generative tasks. Methods like (Yang et al., 2024; Wu et al., 2024b; Li et al., 2024a; Yang et al., 2023; He et al., 2024; Zhang et al., 2025) enhanced Gaussians’ flexibility in handling dynamic and complex scenes by learning 3D Gaussians in canonical space combined with deformation fields. Zhang et al. (2024a) introduces separated intrinsic and dynamic appearance feature to capture the unchanged scene appearance along with dynamic variation like illumination and weather. Additionally, methods like (Qin et al., 2024; Wu et al., 2024a; Shi et al., 2024) incorporate language information into 3DGS scene representation, enabling high-quality open-vocabulary 3D scene queries.
@@
Sparse View Reconstruction. Although techniques of NeRF and 3DGS perform well in dense-view scenarios, they face challenges in sparse-view settings, particularly in maintaining high-quality scene generation under sparse data conditions, necessitating further optimizations to improve generation quality and efficiency. Methods like (Yu et al., 2021; Charatan et al., 2024; Chen et al., 2024a) use a feed-forward manner to learn the 3D scene, which take images as input and predict the parameters of the 3D representation. Specifically, PixelNeRF (Yu et al., 2021) uses convolutional neural networks to extract features from input contexts, addressing the challenges of low-quality reconstructions caused by sparse views. PixelSplat leverages epipolar transformers to extract scene features, while MVSplat uses cost volumes to improve scene feature extraction. FreeNeRF further improves reconstruction by employing frequency and density regularization strategies to reduce artifacts from insufficient inputs, without additional computational costs. To mitigate overfitting in 3DGS (Kerbl et al., 2023) models under sparse views, approaches like FSGS (Zhu et al., 2023) and SparseGS (Xiong et al., 2023) introduce depth estimators to regularize the optimization process. FreeSplat (Wang et al., 2024) builds on this by incorporating not only cost volume construction but also multi-scale feature aggregation, allowing accurate 3D Gaussian distribution localization and supporting free-viewpoint synthesis. These advancements demonstrate the growing potential of sparse-view reconstruction technologies. Learning from the 3DGS-based generative frameworks (Charatan et al., 2024; Chen et al., 2024a; Wang et al., 2024; Zhang et al., 2024b; Xu et al., 2024; Chen et al., 2024b; Ye et al., 2024), SLGaussian adopts a similar approach to infer high-quality 3DGS semantic scenes in sparse-view semantic learning tasks.
@@
3D Semantic Field. Integrating semantic information into 3D scenes is crucial for downstream fields such as robotics, AR, and VR. 2D semantic segmentation mathods like (Kirillov et al., 2023) and 2D open-vocabulary segmentation methods like (Li et al., 2022; Ghiasi et al., 2022; Liang et al., 2023) have inspired further exploration in understanding 3D scenes. Zhi et al. (2021) first introduced semantic layers in NeRF, laying the foundation for deeper scene understanding. Subsequent studies, such as LERF (Kerr et al., 2023), ISRF (Goel et al., 2023), and DFF (Yen-Chen et al., 2022), have incorporated more refined semantic segmentation techniques, significantly improving the level of detail in 3D visual representations. Furthermore, 3DGS technology has also made breakthroughs in semantic tasks. The foundational work of (Kerbl et al., 2023) and the subsequent developments of EgoLifter (Gu et al., 2024), SAGA (Cen et al., 2023), Gaussian Grouping (Ye et al., 2023), and Gaga (Lyu et al., 2024) have greatly optimized the application of 3DGS in semantic segmentation. LangSplat(Qin et al., 2024), LEGaussians(Shi et al., 2024), OpenGaussian(Wu et al., 2024a) and FastLGS (Ji et al., 2024) further explored the integration of language information into 3DGS, deepening semantic interpretation in 3D scene modeling. More previous methods like (Chen et al., 2023; Ding et al., 2023; Ha and Song, 2022; Huang et al., 2023; Jatavallabhula et al., 2023; Kobayashi et al., 2022; Mazur et al., 2023; Peng et al., 2023; Takmaz et al., 2023) have also explored embedding open-vocabulary information into 3D scenes to enable open-vocabulary interaction. However, these techniques often exhibit limitations in handling sparse view inputs, especially when viewpoints are highly limited, making it difficult to effectively manage semantic consistency between views. The lack of sufficient RGB information leads to suboptimal reconstruction quality of both 3D scenes and their semantic content, highlighting the urgent need for efficient and accurate 3D semantic learning under sparse views.
@@
Figure 2. Overview of our approach. Starting with two-view RGB images, we first apply SAM (Kirillov et al., 2023) and a video object tracking model to generate consistent labeled masks (Associated Masks) and construct the language memory bank to store CLIP embeddings (Radford et al., 2021) for each label ID (Section 3.2). The RGB images and associated masks are processed by the Base Gaussians Prediction and Semantic Parameters Prediction modules, which are then combined to generate semantic Gaussians. These are then used to render the 3D semantic field (Section 3.1). Finally, we perform Open-vocabulary Querying through our Multi-view Language Memory Bank (Section 3.2 and Section 3.3), enabling efficient querying of objects within the 3D scene.
3. Proposed Approach
@@
3.1. 3DGS-based Semantic Parameters Prediction
@@
Base Gaussians Prediction. Inferring the semantic field directly from RGB images makes it difficult for the model to fully capture the scene’s geometric structure. A robust semantic field representation requires a deep understanding of the scene’s geometry. To address this, we first apply 3D scene reconstruction methods to provide the model with a foundational grasp of the scene, which then guides the learning of semantic features. As a result, we construct a regular color field using a feed-forward 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) method to derive the base Gaussians:
(1) $\{{c}_{i},{o}_{i},{\mu}_{i},{\delta}_{i}\}=\text{GS-Pred}\left\{\{R_{j}^{3\times H\times W}\}\right\},$
@@
here, ${c}_{i}$, ${o}_{i}$, ${\mu}_{i}$, and ${\delta}_{i}$ represent the color, opacity, mean, and covariance of the $i$-th Gaussian, respectively. The term $R_{j}^{3\times H\times W}$ refers to the RGB image from the $j$-th view, where $H$ and $W$ correspond to the height and width of the image. It is important to note that in this module, we have the flexibility to use any feed-forward model. In our work, we utilize an established method (Chen et al., 2024a) to predict the base Gaussians. These base Gaussian parameters are essential for accurately representing both the scene’s geometry and appearance, directly affecting the quality of semantic field reconstruction. The number of Gaussian splats in this method is determined by the input views and the image resolution, ensuring that each pixel is associated with a corresponding Gaussian point. As a result, the prediction of base Gaussians plays a crucial role in structuring the scene representation. We can either train a complete model ourselves or use pre-trained model weights from existing approaches. Once trained, the feed-forward model in this module is frozen to maintain stability for subsequent semantic field learning.
@@
Semantic Parameters Prediction. Besides predicting the base Gaussian parameters, we also predict the 3DGS semantic parameters in another branch. First, we process the input RGB images from different viewpoints to obtain consistent 2D segmentation maps across views, denoted as $\{S_{i}^{1\times H\times W}\}$, where $H$ and $W$ represent the height and width of the image, respectively. Then, during the construction of the language memory bank, which will be detailed later, we map the segmentation maps into semantic label images $\{M_{i}^{3\times H\times W}\}$. The dimensional extension to 3 is necessary because a 1D index space cannot effectively separate the semantic labels representing different objects within the scene. Next, we apply a simple CNN network to extract features from $M_{i}$. After being input into the network, $M_{i}$ first passes through a convolutional layer and then through multiple downsampling operations based on Residual Blocks (He et al., 2016). Finally, it is upsampled back to the original resolution and added to the original input $M_{i}$. The resulting feature map is flattened and then passed through an MLP, which maps it to the Gaussian semantic feature parameters $f_{i}$. These parameters are 3-dimensional, corresponding directly to the 3D labels of $M_{i}$. Finally, $f_{i}$ is combined with the previously obtained base Gaussian parameters $o_{i},\mu_{i},\delta_{i}$ to generate the final 3D semantic field representation $\{o_{i},\mu_{i},\delta_{i},f_{i}\}$.
@@
(2) $G(x)=\exp\left(-\frac{1}{2}(x-\mu)^{\top}\Sigma^{-1}(x-\mu)\right),$
@@
(3) $F(v)=\sum_{i\in\mathcal{N}}f_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$
here, $F(v)$ represents the rendered semantic feature at pixel $v$ for a specific view, and $\mathcal{N}$ denotes the number of Gaussians in the tile. $\alpha_{i}=o_{i}G_{i}^{2D}(v)$, where $G_{i}^{2D}(\cdot)$ refers to the projection of the $i$-th Gaussian onto the 2D image plane.
3.2. Multi-view Language Memory Bank
Mask Association. Directly using the multi-view segmentation results inferred by the SAM model as input for our model can lead to confusion, as the inconsistent 2D segmentations across views interfere with the spatial mapping of semantic labels, as shown in our supplementary material. Existing methods mitigate this issue by using dense viewpoint information and multi-scale segmentation of SAM results. However, when reduced to the extreme sparse scenario, especially when only two views are available, this multi-scale approach to SAM segmentation becomes ineffective. Research on Gaussian scene segmentation (Ye et al., 2023; Dou et al., 2024) treats multi-view images as consecutive frames in a video and applies a video object tracking technique (Cheng et al., 2023) to unify the multi-view 2D segmentation masks inferred by SAM, achieving relatively good results. We adopted this method to enforce multi-view consistency on SAM’s inference results. However, when only two sparse-view images are directly processed by the video tracking model, the lack of sufficient reference views often result in poor segmentation consistency. To ensure more stable processing results for sparse-view image inputs, we replicated the input images five times to create a 10-frame sequence and processed it using the model, resulting in a set of 10 segmentation images with matched indices: $\{T_{11},T_{12},\dots,T_{15},T_{21},T_{22},\dots,T_{25}\}$, where $T_{ij}$ represents the $j$-th frame segmentation corresponding to the $i$-th view. Each pixel in the segmentation images is labeled with the corresponding object index. For each pixel $v$ in the $i$-th view, we choose the category through a voting algorithm:
(4) $S_{i}(v)=\operatorname*{argmax}_{x\in\{T_{ij}(v)\}}\text{count}(x),$